OpenStudy (anonymous):

Write an equation in slope-intercept form for an equation of a line perpendicular to y = -3x +1 and that intersects at point (1, 2/3).

OpenStudy (fifciol):

\[y=\frac{ 1 }{ 3 }x + \frac{ 1 }{ 3 }\]

OpenStudy (texaschic101):

a perpendicular line will have a negative inverted slope....which is -1/3 y = mx + b using slope(m) 1/3 and y = 2/3 and x = 1 2/3 = 1/3(1) + b 2/3 = 1/3 + b 2/3 - 1/3 = b 1/3 = b y = 1/3x + 1/3 Fifciol is correct :)